In a recent study on world happiness, participants were asked to evaluate their current lives on a scale from 0 to 10, where 0 represents the worst possible life and 10 represents the best possible life. The responses were normally distributed, with a mean of 5.4 and a standard deviation of 2.2. Find the probability that a randomly selected study participant’s response was b. more than 9.
Answer #1

Solution :

Given that ,

mean = = 5.4

standard deviation = = 2.2

P(x > 9) = 1 - p( x< 9)

=1- p P[(x - ) / < (9 - 5.4) / 2.2]

=1- P(z < 1.64)

Using z table,

= 1 - 0.9495

= 0.0505